# BLAST: A Stealthy Backdoor Leverage Attack against Cooperative Multi-Agent Deep Reinforcement Learning based Systems

### v1

@@ V-A4 Evaluation Metrics
We use the metrics in Table I to evaluate the effectiveness and stealthiness of BLAST. These metrics quantify the impact of BLAST on models, where the higher ASR and AER values are, the more effective BLAST is, and the lower CPVR and WRVR values are, the more stealthy BLAST is.
@@ V-B Attack Results in SMAC
We further analyze the behaviors of the BLAST agent and clean agents during the attack period to explore how the BLAST agent affects the clean agents and thus leads to the team’s failure. Fig. 5(a) and 5(b) show the snapshots of episodes of VDN agents in 8m and 2s3z maps, respectively. Firstly, we control an enemy unit to execute the trigger behavior. Once the trigger behavior is fully executed, the backdoor in the BLAST agent is triggered, as shown in Fig. 5(a)-(a) and 5(b)-(a). Then, the BLAST agent begins to perform malicious actions, specifically leaving the main battlefield and moving within the sight of its teammates, successfully inducing some teammates to move to unfavorable positions, resulting in other teammates being attacked by enemy units, as shown in Fig. 5(a)-(b-c) and 5(b)-(b-c). In the end, all teammates of the BLAST agent die, followed by the death of the BLAST agent itself, and the agent team fails, as shown in Fig. 5(a)-(d) and 5(b)-(d).
@@ V-C Attack Results in Pursuit
Figure 7: The average episode rewards of BLAST model of VDN and QMIX.
@@
To verify the universality of our BLAST attacks, we conduct BLAST in the Pursuit environment [30]. Fig. 7 shows the training process of BLAST models of VDN and QMIX. We can see that the average episode reward in poisoned episodes gradually decreases as training progresses, while the average episode reward in clean episodes eventually converges to the original level. This indicates that our BLAST backdoor attacks are effective.
We further analyze the time-step reward of each agent in clean and poisoned episodes under the same random seed, as shown in Fig. 8. Note that the Pursuit environment supports viewing the individual reward value of each agent and the team reward is the average of the reward values of each agent. An increase in the reward of a certain agent indicates that it participates in capturing an evader and successfully captures it. It is not difficult to observe that the rewards of clean agents are not affected before the BLAST attack begins. In the poisoned episode, the reward values of all clean agents during the attack period decrease compared to the clean episode. This indicates that during the attack period, these agents do not collaborate well to capture evaders, that is, the BLAST agent plays a unilateral malicious role. After the attack period, we can find that most agents do not quickly experience an increase in rewards. This is because due to the lag effect of the BLAST attack (i.e., the BLAST agent guides some of the clean agents to the unfavorable position), the agents cannot immediately capture the remaining evaders, but need some time to do so. Once the BLAST attack is performed successfully, the agent team will spend more time capturing all the evaders.
@@ V-D Resistance to Backdoor Defense Methods
To further validate the stealthiness of our backdoors, we evaluate whether existing backdoor defense methods can detect our backdoors in VDN. We use two popular defense methods (activation clustering [18] and spectral signature [19]) to conduct this evaluation. We use the BLAST model to interact with the 8m environment and sample 28000 clean samples and 950 poisonous samples according to action types, dividing them into six sets (including 5 movement actions and 1 attack action). Following activation clustering, we extract the activations in the penultimate layer of the model and use principal component analysis to downscale the activations to 3 dimensions and then cluster them using k-means with $k=2$. The clustering analysis results are shown in Fig. 9, in which we can discover the activations of poisonous samples are not distinguishable from the clean ones, and the prediction results are basically incorrect. Besides, the detection results of spectral signature are shown in Figure 9, where we can find that the distribution of outlier scores for the poisonous samples is roughly the same as that for the clean samples, with no significant deviation.
@@ V-E1 $\lambda$ in Hacked Reward
We further evaluate the performance of our BLAST backdoor attacks against VDN with different values of $\lambda\in\{0,0.25,0.50,0.75,1\}$ in hacked reward function, as shown in Fig 11 (a). We can find that when $\lambda=0$, the $ASR$ can reach up to 84.4%. This indicates that including only the first term in the hacked reward, i.e., target failure state guiding is useful to some extent against c-MADRL. However, when $\lambda=1$, the $ASR$ is only 16.4%, although it can maintain a very low $CPVR$ which is 1.1%. This suggests that the attack is not very effective when considering only the second term in the hacked reward, i.e., the effect of the BLAST agent on the next time-step actions of its teammates. Attacks perform best when $\lambda=0.5$, being able to maintain a 1.6% $CPVR$ and achieve a 96.7% $ASR$. This further demonstrates the effectiveness of our hacked reward for backdoor attacks against c-MADRL.
@@ V-E2 Poisoning Rate $pr$
We further evaluate the performance of our BLAST backdoor attacks with different values of poisoning rate $pr\in\{2\%,5\%,10\%,20\%\}$, as shown in Fig 11 (b). We can observe that as the poisoning rate increases, $ASR$ continues to rise, but $CPVR$ also continues to increase. When the poisoning rate is 5%, $ASR$ and $CPVR$ are a good balance point, which can make the attack effective while ensuring good concealment.
@@ VI-A Backdoor Attack against DRL
Most existing studies on backdoor attacks focuse on DNN [33]. Due to the representation ability and feature extraction ability inherited from DNN, DRL also inevitably faces this threat. Kiourti et al. [5] propose a backdoor attack method against DRL with image patch triggers. During training, when the model’s input contains a trigger and the output is the target action or a random one, they maximize the corresponding reward value to achieve a targeted or untargeted attack. Yang et al. [6] investigate the persistent impact of backdoor attacks against LSTM-based PPO algorithm [29]. When an image patch trigger appears at a certain time step, the backdoor agent persistently goes for the attacker-specified goal instead of the original goal. Ashcraft et al. [8] use in-distribution triggers and multitask learning to train a backdoor agent. An in-distribution trigger is an observation that is not anomalous to the environment and is therefore more difficult to detect. Wang et al. [7] study backdoor attacks against DRL in two-player competitive games. The opponent’s actions are used as a trigger to switch the backdoor agent to fail fast. Wang et al. [9] study backdoor attacks on DRL-based traffic congestion control systems and used a specific set of combinations of positions and speeds of vehicles in the observation as the trigger. When the trigger is present, the backdoor DRL model could generate malicious decelerations to cause a physical crash or traffic congestion. Cui et al. [10] propose a sparse backdoor attack that selects observations with high attack values to inject triggers and generates trigger patterns using mutual information-based tuning. Chen et al. [11] study backdoor attacks against offline reinforcement learning, which poison open-source reinforcement learning datasets to implant backdoors to DRL agents. They use instant triggers and explore the effectiveness of attacks with distributed strategy and one-time strategy. Our previous works [32, 25] studied a new temporal and spatial-temporal backdoor trigger to DRL, which has shown high concealment.
@@ VI-B Backdoor Attack against c-MADRL
So far, only a few studies have been done on backdoors in c-MADRL. Chen et al. [13] first study backdoor attacks against c-MADRL. They consider both out-of-distribution and in-distribution triggers and use a pre-trained expert model to guide the selection of actions of poisoned agents. Besides, they modify the team reward to encourage poisoned agents to exhibit their worst actions. Their method implants backdoors in all agents, and every agent that observes the trigger will perform an instant malicious action. Chen et al. [14] propose backdoor attacks that affect the entire multi-agent team by targeting just one backdoor agent. During training, they introduce a random network distillation module and a trigger policy network to guide the backdoor agent on what actions to select and when to trigger the backdoor, thereby enabling the agent to adversely affect other normal agents. During execution, the backdoor agent needs to rely on the trigger policy network to determine whether to trigger the backdoor. Zheng et al. [15] also implant a backdoor into one of the agents in a multi-agent team, and use distance as a condition to trigger the backdoor. When the trigger condition is met, the backdoor agent performs a specified action which leads to the failure of the team task. These existing backdoor attacks are effective against c-MADRL, but they either poison all agents, require additional networks to activate backdoors, or do not take into account the influence between agents. We compare these existing backdoor attacks against DRL in Table III.
@@ VI-C Backdoor Defense Mechanism
Existing backdoor defenses broadly include backdoor detection, data filtering, and backdoor mitigation. [19, 18] are two poisoned input detection methods that depend on the difference of covariance spectrum of the feature representation and activation value distributions in the last hidden neural layer between poisonous and clean samples, respectively. Wang et al. [34] use reverse engineering to infer and detect the shapes and locations of backdoor triggers and mitigate backdoors through input filters, neuron pruning, and unlearning techniques. Neuralsanitizer [35] is a backdoor detection and removal tool that identifies real triggers and detects backdoors by reconstructing a set of potential triggers and removing redundant features. For backdoor attacks against DRL, PolicyCleanse [36] is a backdoor detection and mitigation approach that uses policy optimization with a reversed cumulative reward of the backdoor agent in competitive DRL policies. BIRD [37] is a method for detecting and removing backdoors in DRL policies by analyzing the attack process and unique behaviors of the backdoor agent, formulating the trigger restoration as an optimization problem, and further mitigating the backdoor through fine-tuning. Sanyam et al. [38] develop a classifier that utilizes the neuron activation patterns in DRL to detect the presence of in-distribution backdoor triggers. Since BLAST is hidden in only one of the multiple DRL agents whose triggers and actions are decoupled and distributed in different time series, rather than in individual input, it is difficult to override the trigger pattern by means of fine-tuning. More sophisticated defense methods need to be developed.
@@ VII Conclusions And Future
In this paper, we study backdoors against c-MADRL. To enhance the stealthiness, effectiveness, and practicality of backdoor attacks, we propose a novel backdoor attack against c-MADRL, BLAST, which can disrupt the entire multi-agent team by implanting a backdoor in only one agent. We use spatiotemporal features rather than an instant state as the backdoor trigger and design the malicious reward function based on unilateral influence. We evaluate our BLAST against VDN, QMIX, and MAPPO algorithms in 2 c-MADRL environments and 2 backdoor defense methods. The experimental results demonstrate that BLAST can achieve a great attack success rate and a low clean performance variance rate. In the future, we will explore backdoor attacks in black-box scenarios, as well as study effective defense methods for c-MADRL backdoors.

### v2

@@ V-A4 Evaluation Metrics
We use the metrics in Table I to evaluate the effectiveness and stealthiness of BLAST. These metrics quantify the impact of BLAST on DRL policies, where the higher ASR and AER values are, the more effective BLAST is, and the lower CPVR and WRVR values are, the more stealthy BLAST is.
@@ V-B Attack Results in SMAC
We further analyze the behaviors of the BLAST agent and clean agents during the attack period to explore how the BLAST agent affects the clean agents and thus leads to the team’s failure. Fig. 5 shows the snapshots of episodes of VDN agents in 8m and 2s3z maps, respectively. Firstly, we control an enemy unit to execute the trigger behavior. Once the trigger behavior is fully executed, the backdoor in the BLAST agent is triggered, as shown in Fig. 5(a)-(a) and 5(b)-(a). Then, the BLAST agent begins to perform malicious actions, specifically leaving the main battlefield and moving within the sight of its teammates, successfully inducing some teammates to move to unfavorable positions, resulting in other teammates being attacked by enemy units, as shown in Fig. 5(a)-(b-c) and 5(b)-(b-c). In the end, all teammates of the BLAST agent die, followed by the death of the BLAST agent itself, and the agent team fails, as shown in Fig. 5(a)-(d) and 5(b)-(d).
@@ V-C Attack Results in Pursuit
Figure 7: The average episode rewards of the BLAST model of VDN and QMIX.
@@
To verify the universality of our BLAST attacks, we conduct BLAST in the Pursuit environment [33]. Fig. 7 shows the training process of BLAST policies of VDN and QMIX. We can see that the average episode reward in poisoned episodes gradually decreases as training progresses, while the average episode reward in clean episodes eventually converges to the original level. This indicates that our BLAST backdoor attacks are effective.
We further analyze the time-step reward of each agent in clean and poisoned episodes under the same random seed in Fig. 8. Note that the Pursuit environment supports viewing the individual reward value of each agent and the team reward is the average of the reward values of each agent. An increase in the reward of a certain agent indicates that it participates in capturing an evader and successfully captures it. It is not difficult to observe that the rewards of clean agents are not affected before the BLAST attack begins. In the poisoned episode, the reward values of all clean agents during the attack period decrease compared to the clean episode. This indicates that during the attack period, these agents do not collaborate well to capture evaders; that is, the BLAST agent plays a unilateral malicious role. After the attack period, we can find that most agents do not quickly experience an increase in rewards. This is because due to the lag effect of the BLAST attack (i.e., the BLAST agent guides some of the clean agents to the unfavorable position), the agents cannot immediately capture the remaining evaders, but need some time to do so. Once the BLAST attack is performed successfully, the agent team will spend more time capturing all the evaders.
@@ V-D Resistance to Backdoor Defense Methods
To validate the stealthiness of our BLAST attack, we investigate its detectability against existing defense mechanisms in VDN. Three established approaches are evaluated: activation clustering [18], spectral signature analysis [19] (both designed for DNNs), and BIRD [35] (developed for DRL). We use the BLAST policy to interact with the 8m environment and sample 28000 clean samples and 950 poisonous samples according to action types, dividing them into six sets (including 5 movement actions and 1 attack action). Following activation clustering, we extract the activations in the penultimate layer of the model and use principal component analysis to downscale the activations to 3 dimensions and then cluster them using k-means with $k=2$. The clustering analysis results are shown in Fig. 9, in which we can discover that the activations of poisonous samples are not distinguishable from the clean ones, and the prediction results are incorrect. Besides, the detection results of spectral signature are shown in Figure 9, where we can find that the distribution of outlier scores for the poisonous samples is roughly the same as that for the clean samples, with no significant deviation.
We further employ BIRD [35] to evaluate VDN, QMIX, and MAPPO models injected with BLAST across distinct maps (“8m”, “3m”, and “2s3z”) in the StarCraft2Env. Three maps are configured with specific triggers to generate backdoor models: $\mathcal{T}_{1}$ is deployed on “8m” and “2s3z”, while $\mathcal{T}_{2}$ is applied to “3m”. For each backdoor model, we use BIRD with an initial perturbation magnitude of 0.1 to generate a “pseudo-trigger”. Then, these backdoor models are run to interact with the environment with and without pseudo-triggers, respectively. BIRD will use the maximum consecutive reward drop plus its standard deviation during those interactions without pseudo-triggers as the backdoor detection threshold to detect backdoor attacks. The detection results show that in the “8m” map, BIRD fails to detect backdoor attacks under the VDN, QMIX, and MAPPO algorithms; in the “3m” map, BIRD only succeeds under VDN, but still fails under QMIX and MAPPO; and in the “2s3z” map, BIRD only succeeds under MAPPO, but fails under VDN and QMIX. The core method of BIRD’s backdoor detection is to compare the abnormally large reward drop during the pseudo-trigger occurs with the reward changes when there is no pseudo-trigger. Since our BLAST attack delays the malicious action generation until after the backdoor is triggered, rather than being generated during the trigger occurs like existing attacks, the reward does not exhibit anomalous changes in the presence of pseudo triggers.
@@ V-E1 $\lambda$ in Hacked Reward
We further evaluate the performance of our BLAST backdoor attacks against VDN with different values of $\lambda\in\{0,0.25,0.50,0.75,1\}$ in hacked reward function, as shown in Fig 11 (a). We can find that when $\lambda=0$, the $ASR$ can reach up to 84.4%. This indicates that including only the first term in the hacked reward, i.e., target failure state guiding, is useful to some extent against c-MADRL. However, when $\lambda=1$, the $ASR$ is only 16.4%, although it can maintain a very low $CPVR$, which is 1.1%. This suggests that the attack is not very effective when considering only the second term in the hacked reward, i.e., the effect of the BLAST agent on the next time-step actions of its teammates. Attacks perform best when $\lambda=0.5$, being able to maintain a 1.6% $CPVR$ and achieve a 96.7% $ASR$. This further demonstrates the effectiveness of our hacked reward for backdoor attacks against c-MADRL.
@@ V-E2 Poisoning Rate $pr$
We further evaluate the performance of our BLAST backdoor attacks with different values of poisoning rate $pr\in\{2\%,5\%,10\%,20\%\}$, as shown in Fig. 11 (b). We can observe that as the poisoning rate increases, $ASR$ continues to rise, but $CPVR$ also continues to increase. When the poisoning rate is 5%, $ASR$ and $CPVR$ are a good balance point, which can make the attack effective while ensuring good concealment.
@@ VI-A Backdoor Attack against DRL
Most existing studies on backdoor attacks focus on DNN [38]. Due to the representation ability and feature extraction ability inherited from DNN, DRL also inevitably faces this threat. Kiourti et al. [36] propose a backdoor attack method against DRL with image patch triggers. During training, when the model’s input contains a trigger and the output is the target action or a random one, they maximize the corresponding reward value to achieve a targeted or untargeted attack. Yang et al. [6] investigate the persistent impact of backdoor attacks against the LSTM-based PPO algorithm [29]. When an image patch trigger appears at a certain time step, the backdoor agent persistently goes for the attacker-specified goal instead of the original goal. Ashcraft et al. [8] use in-distribution triggers and multitask learning to train a backdoor agent. An in-distribution trigger is an observation that is not anomalous to the environment and is therefore more difficult to detect. Wang et al. [7] study backdoor attacks against DRL in two-player competitive games. The opponent’s actions are used as a trigger to switch the backdoor agent to fail fast. Wang et al. [9] study backdoor attacks on DRL-based traffic congestion control systems and used a specific set of combinations of positions and speeds of vehicles in the observation as the trigger. When the trigger is present, the backdoor DRL model could generate malicious decelerations to cause a physical crash or traffic congestion. Cui et al. [32] propose a sparse backdoor attack that selects observations with high attack values to inject triggers and generates trigger patterns using mutual information-based tuning. Chen et al. [10] study backdoor attacks against offline reinforcement learning, which poison open-source reinforcement learning datasets to implant backdoors to DRL agents. They use instant triggers and explore the effectiveness of attacks with a distributed strategy and a one-time strategy. Our previous works [37, 5] studied a new temporal and spatial-temporal backdoor trigger to DRL, which has shown high concealment.
@@ VI-B Backdoor Attack against c-MADRL
So far, only a few studies have been done on backdoors in c-MADRL. Chen et al. [13] study backdoor attacks against c-MADRL. They consider both out-of-distribution and in-distribution triggers and use a pre-trained expert model to guide the selection of actions of poisoned agents. Besides, they modify the team reward to encourage poisoned agents to exhibit their worst actions. This method implants backdoors in all agents, and all agents will perform an instant malicious action when observing the trigger. Chen et al. [14] propose a backdoor attack that affects the entire multi-agent team by targeting just one backdoor agent. This attack requires a random network distillation module and a trigger policy network to guide the backdoor agent on what actions to select and when to trigger the backdoor. During execution, the backdoor agent needs to rely on the trigger policy network to determine whether to trigger the backdoor. Zheng et al. [15] also implant a backdoor into one of the agents in a multi-agent team, and use distance as a condition to trigger the backdoor. When the trigger condition is met, the backdoor agent performs a specified action which leads to the failure of the team task. These existing backdoor attacks are effective against c-MADRL, but they either poison all agents, require additional networks to activate backdoors, or do not take into account the influence between agents. We compare these existing backdoor attacks against DRL in Table III.
@@ VI-C Backdoor Defense Mechanism
Existing backdoor defenses broadly include backdoor detection, data filtering, and backdoor mitigation. [19, 18] are two poisoned input detection methods that depend on the difference of covariance spectrum of the feature representation and activation value distributions in the last hidden neural layer between poisonous and clean samples, respectively. Wang et al. [39] use reverse engineering to infer and detect the shapes and locations of backdoor triggers and mitigate backdoors through input filters, neuron pruning, and unlearning techniques. Neuralsanitizer [40] is a backdoor detection and removal tool that identifies real triggers and detects backdoors by reconstructing a set of potential triggers and removing redundant features. For backdoor attacks against DRL, PolicyCleanse [41] is a backdoor detection and mitigation approach that uses policy optimization with a reversed cumulative reward of the backdoor agent in competitive DRL policies. Sanyam et al. [42] develop a classifier that utilizes the neuron activation patterns in DRL to detect the presence of in-distribution backdoor triggers. Both BIRD [35] and SHINE[43] claim to detect backdoor policies both in single- and multi-agent DRL scenarios. BIRD requires accessing the clean and poisoned environments to compare reward change behaviors and restore the potential trigger. SHINE collects a set of poisoned trajectories to identify features in the trigger-presenting states leading to backdoor actions and deems them as the backdoor trigger. All existing defense methods assume that backdoor actions are immediately generated when the trigger occurs. Since BLAST is hidden in only one of multiple DRL agents, whose triggers and actions are decoupled and distributed in different time series, rather than in individual input, most of the existing methods will fail to detect this attack-trigger-decoupled attack.
@@ VII Conclusions and Future
In this paper, we study backdoors against c-MADRL. To enhance the stealthiness, effectiveness, and practicality of backdoor attacks, we propose a novel backdoor attack against c-MADRL, BLAST, which can disrupt the entire multi-agent team by implanting a backdoor in only one agent. We use spatiotemporal features rather than an instant state as the backdoor trigger and design the malicious reward function based on unilateral influence. We evaluate our BLAST against VDN, QMIX, and MAPPO algorithms in two c-MADRL environments and three backdoor defense methods. The experimental results demonstrate that BLAST can achieve a high attack success rate and a low clean performance variance rate. In the future, we will explore backdoor attacks in black-box scenarios, as well as study effective defense methods for c-MADRL backdoors.